OpenStudy (anonymous):

the sum of three consecutive even integers is -450 find the integers

OpenStudy (anonymous):

Same idea as last time. a + b + c = -450, and a, b, and c are consecutive integers, but they are all even. That means b=a+2 and c=b+2. That implies c=(a+2)+2, or c=a+4. a + a + 2 + a + 4 = -450 Combine like terms. 3a + 6 = -450 Use algebra. 3a+6-6=-450-6 3a=-456 3a/3=-456/3 a=-152 You have your answer for the first integer. Now, since b=a+2 and c=a+4, you just have to add. b=a+2 b=-152+2 b=-150 c=a+4 c=-152+4 c=-148 -152, -150, -148.